OpenStudy (anonymous):

Can someone please help me find the sum of this sigma notation? \[\sum_{k=1}^{10}2k^3\]

OpenStudy (jtvatsim):

Alright, this one is not too bad... Do you know how to interpret the sigma notation?

OpenStudy (anonymous):

no :(

OpenStudy (jtvatsim):

OK, here's a quick review of how to read sigma notation. Let's use an easy first question. \[\sum_{k=1}^{3} k\]

OpenStudy (jtvatsim):

So what the weird sigma thing means is to add up a bunch of numbers.

OpenStudy (jtvatsim):

the bottom part (k=1) tells you what to plug in first, and the top part (3) tells you when to stop.

OpenStudy (jtvatsim):

So, we just get \[\sum_{k=1}^3 k = (1) + (2) + (3) = 6\]

OpenStudy (anonymous):

so keep plugging in numbers until you get 3?

OpenStudy (anonymous):

OH i see!!

OpenStudy (jtvatsim):

does that make sense?

OpenStudy (jtvatsim):

great! so for the question you were trying to answer....

OpenStudy (jtvatsim):

the first thing is going to be 2(1)^3

OpenStudy (jtvatsim):

then, we need to add 2(2)^3...

OpenStudy (jtvatsim):

then, 2(3)^3 and so on

OpenStudy (jtvatsim):

it's a bit long, but not too hard... :)

OpenStudy (anonymous):

I gotcha! thank you SO much for taking the time to explain everything to me, I really appreciate it! :)

OpenStudy (jtvatsim):

no problem!

OpenStudy (anonymous):

so to find the "sum" when I get 2(1)^3...2(10)^3 do i add all of them up?

OpenStudy (jtvatsim):

that's right, you should get a pretty big number... if I did the adding right, you should get 6050

OpenStudy (anonymous):

thank you! :)

OpenStudy (jtvatsim):

yw

OpenStudy (anonymous):

@jtvatsim for this problem instead of starting at 1 would i start at 0?? \[\sum_{j=0}^{4}(j^2+1)\] so it would be (0)^2+1, (1)^2 +1, (2)^2+1, (3)^2+1 and (4)^2 +1?

OpenStudy (jtvatsim):

you've got it! You're on a roll now. :)

OpenStudy (jtvatsim):

Sigma notation just looks big and scary, but it's really actually nice. lol

OpenStudy (anonymous):

thanks so much! no kidding :p

OpenStudy (anonymous):

so for this one \[\sum_{k=1}^{\infty}8(\frac{1}{10})^k\] how would i solve this??

OpenStudy (jtvatsim):

OK, this one is a bit different since using the method we've been doing would require us to find an infinite amount of numbers (no thank you...)

OpenStudy (jtvatsim):

Instead, let's try to get an idea of what is going on by writing out just a few terms... \[\sum_{k=1}^\infty 8(1/10)^k = 8(1/10) + 8(1/10)^2 + 8(1/10)^3 + ...\]

OpenStudy (jtvatsim):

of course, we have to keep going forever, but notice that we can factor out the 8 showing us that we really have \[8[ (1/10) + (1/10)^2 + (1/10)^3 + ... ]\]

OpenStudy (jtvatsim):

OK, great, but now what??? Well, notice that this is a geometric series since we are just multiplying by (1/10) each time.

OpenStudy (jtvatsim):

The geometric series formula is \[\frac{ a }{ 1-r }\] where a is the first term and r is the ratio.

OpenStudy (anonymous):

so even though we know it is multiplying by 1/10 each time is there a way to "figure out" infinity?

OpenStudy (jtvatsim):

Yes, that's what the geometric series formula does... it adds up infinitely many things that multiply by some ratio each time.

OpenStudy (jtvatsim):

So, it turns out (by the geometric series formula) that \[[(1/10) + (1/10)^2 + (1/10)^3 + ...] = \frac{ 1/10 }{ 1-(1/10) } = 1/9\]

OpenStudy (jtvatsim):

Now, remember that we still had the 8 out front so the final answer is 8*(1/9) = 8/9

OpenStudy (anonymous):

so 8/9 would be infinity? in this equation at least?

OpenStudy (jtvatsim):

Hmm, not quite. What the sigma was asking us was to start with k=1 and keep adding FOREVER! You might think that if you add forever you would never get a "final" answer. However, math is a bit surprising, so what we have found (using the tricky geometric series) is that if we were to add these numbers FOREVER (the infinity) we would actually get the answer 8/9. Does that make sense? Infinity tends to be hard to imagine. :)

OpenStudy (anonymous):

oh okay, that makes more sense to me now :P you're awesome! i wish i could give you an infinite amount of medals :D thanks again :)

OpenStudy (jtvatsim):

LOL, that would be nice, but that might take up all the memory on the internet! :D

OpenStudy (anonymous):

hahaha!